# Multi-View Video-Based Learning: Leveraging Weak Labels for Frame-Level Perception

### 18 Mar 2024 v1

@@ 2 Literature Review
The literature in multi-view learning focuses on efficiently utilizing data from different sources for various perception tasks [1, 2, 3]. A detailed survey of the multi-view learning literature are presented by Wu et al. [6], Zhou et al. [1] and Luo et al. [2]. In literature, deep learning-based multi learning [2, 3, 7, 8] report enhanced detection accuracy. Su et al. [9] introduce a multi-view CNN architecture for 3D shape recognition, and show that using multiple views enhances the recognition accuracy. In recent years, transformer-based multi-learning report state-of-the-art accuracy [5] for multi-view learning. Apart from deep learning, other techniques such as conditional random field (CRF) [10], discriminant analysis [11], and point-wise mutual information [12] are used for multi-view learning.
While existing work has made significant contributions to multi-view learning, the problem of using weak labels for frame-level perception in multi-view learning has not been addressed in detail. Amongst the literature, Yasuda et al. [5] propose a transformer framework which is trained using bags of labels for a sequence, as weak labels, for event detection. The trained transformer are then used for frame-level event detection, by extracting the outputs from the latter layers.

### 19 Mar 2024 v2

@@ 2 Literature Review
The literature in multi-view learning focuses on efficiently utilizing data from different sources for various perception tasks [1, 2, 3]. A detailed survey of the multi-view learning literature are presented by Yan et al. [1], Li et al. [2] and Xu et al. [3]. In literature, deep learning-based multi-view learning [6, 7, 1, 8, 9] report enhanced detection accuracy. Su et al. [7] introduce a multi-view CNN architecture for 3D shape recognition, and show that using multiple views enhances the recognition accuracy. In recent years, transformer-based multi-learning report state-of-the-art accuracy [5] for multi-view learning. Apart from deep learning, other techniques such as conditional random field (CRF) [10] and discriminant analysis [11] are used for multi-view learning.
While existing work has made significant contributions to multi-view learning, the problem of using weak labels for frame-level perception in multi-view learning has not been addressed in detail [12, 13]. Amongst the literature, Yasuda et al. [5] propose a transformer framework which is trained using bags of labels for a sequence, as weak labels, for event detection. The trained transformer are then used for frame-level event detection, by extracting the outputs from the latter layers.